OpenStudy (twalt13):

I dont understand this question how do you start it off ? Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 8, -14, and 3 + 9i

OpenStudy (sirm3d):

multiply the factors (c-8), (x+14), (x-[3+9i]) and (x+[3+9i)

OpenStudy (sirm3d):

the last factor should be (x-[3-9i])

OpenStudy (sirm3d):

and the first factor is (x-8), not (c-8)

OpenStudy (twalt13):

how do you mutiply the (x-[3+9]) (x-[3-9i])

OpenStudy (sirm3d):

regroup them as \[\left( \left[ x-3 \right]+9i \right)\left( \left[ x-3 \right]-9i \right)\]

OpenStudy (twalt13):

then you foil with (x-8) ?

OpenStudy (sirm3d):

FOIL works for two binomials only. i suggest multiply them vertically

OpenStudy (twalt13):

you lost me when u said that example ?

OpenStudy (sirm3d):

i assumed foil means foil method of multiplying binomials. the product \[(x-3+9i)(x-3-9i)\] is a trinomial \[x^2 -6x+90\]

OpenStudy (twalt13):

ok then combine like terms once i do the other side by multipying ?

OpenStudy (sirm3d):

right